# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v1

@@ 2. Estimation procedure:
  10. Output:
    Estimation $\hat{\psi}$ or $\hat{\Psi}$.
  11. Properties:
    Unbiased, with relative MSE (OR) or MSE (LOR) less than $A$;
    Average sample sizes: (149), (150); approximate ratio $\lambda$;
    Efficiency with separate sampling: (152); approaches $1$ for $A$ small
    Average number of groups: (154), (155);
    Efficiency with group sampling: (156).
The relative MSE obtained from simulations is compared with the target $A$ in Figure 15. The simulation consists, as with previous estimators, of $10^{6}$ realizations for each combination of parameters. The difference between simulation results and target is seen to increase with $A$ and with $\phi$, as for RR. However, in OR the difference observed for large $\phi$ vanishes when $A$ is small, unlike in RR (compare the leftmost parts of Figures 6(a) and 15(a)). This is related to the fact that the terms $1-\bar{p}_{1}/(s_{1}-2+2p_{1})$ and $1-\bar{p}_{2}/(s_{2}-2p_{2})$ in (130), which are replaced by $1$ in the uniform bound (131), approach $1$ for large $s_{1}$, $s_{2}$. Thus when $A$ is small, which gives large values of $s_{1}$ and $s_{2}$, the uniform bound (131) is almost as good as (130). In RR, on the other hand, the corresponding factors are $(s_{1}-2)(1-p_{1})/(s_{1}-2+2p_{1})<1-p_{1}$ and $1-p_{2}$, as it is seen comparing (27) and (28), and these do not tend to $1$ for large $s_{1}$, $s_{2}$. (In LRR the factors are more cumbersome, see Mendo [16, theorem 2]; but for large $s_{1}$, $s_{2}$ they tend to the same values $1-p_{1}$, $1-p_{2}$ as in RR.)
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.025$, $p_{2}=0.4$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.04$, $p_{2}=0.0025$)
Dashed line: target. Dots: simulation.
Figure 15: Relative MSE for OR with separate sampling
(a) $\lambda=3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
(b) $\lambda=1/3$, $\theta=1/16$, $\phi=0.01$ ($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 16: Normalized average sample sizes for OR with separate sampling
The efficiency of the OR estimator with separate sampling results from particularizing (90) to $\zeta=\psi$, with $\partial\psi/\partial p_{1}=\psi/\bar{p}_{1}$, $\partial\psi/\partial p_{2}=-\psi/\bar{p}_{2}$:
$\eta_{\mathrm{sep}}=\frac{\displaystyle\frac{1}{\operatorname{E}[M_{1}+N_{1}]\bar{p}_{1}}+\frac{1}{\operatorname{E}[M_{2}+N_{2}]\bar{p}_{2}}}{\operatorname{Var}[\psi]/\psi^{2}}.$ (151)
$\eta_{\mathrm{sep}}>\frac{1}{1+A(3r_{1}/2+\mu_{1}+1)},$ (152)
where $\mu_{1}=2$. This is similar to the bound (92) for RR and LRR, the only differences being the $3/2$ factor in (152), arising from the Bernoulli factory, and the last factor in (92) (but note that the latter is approximately $1$ for small $\phi$). By the same arguments used for RR and LRR, $\eta_{\mathrm{sep}}$ for OR approaches $1$ when $A$ tends to 00.
Figure 17 shows simulation results for $\eta_{\mathrm{sep}}$, and compares them with the bound (152). In the simulation, $\eta_{\mathrm{sep}}$ is computed as in previous sections, i.e. using sample averages in (151). The values are seen to be similar to those for RR and LRR, except that with large $A$ the efficiency is slightly lower for OR. Again, this is a consequence of the $3/2$ factor. As in RR and LRR, the difference between simulation and bound tends to be larger when $\phi$ is increased. However, in OR that difference vanishes for small $A$, unlike in the other cases (compare Figure 17(a) with Figure 9(b) or 13(a)). This agrees with the behavior of the MSE discussed earlier.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 17: Efficiency for OR with separate sampling
4.2 Group sampling. Estimation procedure and results
Group sampling for OR estimation consumes a number of groups $G$ given by (94), as with previous estimators, in order to provide the required amounts of samples $M_{i}+N_{i}$ of each population $i=1,2$; and $\operatorname{E}[G]$ can be approximately computed from (96). It is shown in Appendix A.2 that, for $\bar{\phi}$ small, the term $\operatorname{E}[|\Delta|]$ in (96) can be expressed as
$\operatorname{E}\left[|\Delta|\right]\approx\operatorname{E}\left[\left|\frac{3r_{1}/2+s_{1}}{l_{1}p_{1}}-\frac{3r_{2}/2+s_{2}}{l_{2}p_{2}}\right|\right].$ (153)
Thus, proceeding as in Subsection 2.6 and taking into account (142),
where
$Y_{0}=\frac{r_{1}}{2\lambda\bar{\theta}(r_{1}-1)}\left(\lambda\bar{\theta}-1+\sqrt{(\lambda\bar{\theta}-1)^{2}+\frac{4\lambda\bar{\theta}(r_{1}-1)^{2}}{r_{1}^{2}}}\right)$ (155)
and $\mu_{1}=2$. The simulation results for $\operatorname{E}[G]\bar{\phi}$, not shown, are similar to those for $\operatorname{E}[G]\phi$ in RR and LRR, with the difference that for large $A$ the values are slightly greater in OR compared with those two cases, as it has been observed with separate sampling.
where $\operatorname{E}[G]\bar{\phi}$ is given by (154) and (155). Figure 18 represents the theoretical approximation (156), as well as results from simulation. As with separate sampling, the efficiency for large $A$ is slightly lower than in RR and LRR (Figures 11 and 14), and for small $A$ there is almost no difference between simulation results and theoretical approximation even if $\bar{\phi}$ is large.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 18: Efficiency for OR with group sampling
5 Estimation of log odds ratio
5.1 Separate sampling. Estimation procedure and results
The estimation method for the LOR, $\Psi=\log(p_{1}(1-p_{2})/(p_{2}(1-p_{1})))$, has few differences compared to that presented for OR in Section 4. The second stage estimates $\log(p_{1}/(1-p_{1}))$ and $\log(p_{2}/(1-p_{2}))$ separately, using for each the method described in Mendo [16, section 3], which consists of two IBS procedures with the same parameter. Specifically, to estimate $\log(p_{i}/(1-p_{i}))$, $i=1,2$, IBS is applied until $s_{i}$ successes are obtained, which requires a random number $N_{i}^{\prime}$ of observations from population $i$, and then IBS is applied until $s_{i}$ failures are obtained, which requires $N_{i}^{\prime\prime}$ additional observations. These are the same steps as for OR, but with $\alpha=0$. Then, conditioned on $s_{1}$ and $s_{2}$, which are obtained in the first stage, $-H_{N_{i}^{\prime}-1}+H_{N_{i}^{\prime\prime}-1}$ is a conditionally unbiased estimator of $\log(p_{i}/(1-p_{i}))$ with conditional variance less than $1/(s_{i}-5/4)$ for any $p_{i}\in(0,1)$ [16, theorem 3]. Therefore,
is a conditionally unbiased estimator of $\Psi$ with
$\mu_{1}=5/4,\qquad\mu_{2}=5/4,\qquad\mu_{12}=0,$ (159)
The number of samples from population $i=1,2$ used in the second stage, $N_{i}=N_{i}^{\prime}+N_{i}^{\prime\prime}$, has an average
$\operatorname{E}[N_{i}]=\frac{s_{i}}{p_{i}}+\frac{s_{i}}{1-p_{i}}=\frac{s_{i}}{\bar{p}_{i}},$ (160)
where $\bar{p}_{i}$ is defined by (125). Thus, for the same reason as in OR estimation, the first stage for LOR estimation must be based on samples with parameters $\bar{p}_{i}$, $i=1,2$, generated from observations with parameters $p_{i}$. For each $i=1,2$, given $r_{i}$, a random number $\bar{M}_{i}$ of samples with parameter $\bar{p}_{i}$ are generated until $r_{i}$ successes are obtained, which in turn requires $M_{i}$ observations from population $i$. Then $X$ is computed from $\bar{M}_{1}$ and $\bar{M}_{2}$ as in (128). Since (142) holds, imposing the additional condition (143), $\gamma$ and $\delta_{1}$ are obtained from (144) and (145) as for OR, and (22) is satisfied. Once $X$, $\gamma$ and $\delta_{1}$ are known, $s_{1}$ and $s_{2}$ are computed using (64)–(66), as for the other estimators, and then rounded with the restriction (31). The resulting average numbers of samples are bounded by (149) and (150).
$A=\frac{1}{(3r_{1}/2+\mu_{1}+\xi)(r_{1}-1)}.$ (161)
The LOR estimator uses $\xi=1$ and $r_{1}$ given by (81), which in this case is written as
The estimation procedure for LOR is summarized in Algorithm 2, which also lists the properties of the estimator.
Particularizing (90) for $\zeta=\Psi$, with $\partial\Psi/\partial p_{1}=1/\bar{p}_{1}$, $\partial\Psi/\partial p_{2}=-1/\bar{p}_{2}$, and using (149), (150) and the fact that $\operatorname{Var}[\hat{\Psi}]<A$, it follows that the efficiency with separate sampling is bounded by (152) with $\mu_{1}=5/4$. Figure 19 shows the results, which are very similar to those for OR.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: bound. Dots: simulation.
Figure 19: Efficiency for LOR with separate sampling
5.2 Group sampling. Estimation procedure and results
The group sampling procedure is analogous to that described for the other estimators. The average number of groups $\operatorname{E}[G]$ is computed from (96), where $\operatorname{E}[|\Delta|]$ satisfies (153), as shown in Appendix A.2. Thus $\operatorname{E}[G]$ is approximately given by (154) and (155), as for OR, but with $\mu_{1}=5/4$ and with a possibly different computed value for $r_{1}$. Simulation results are omitted.
(a) $\lambda=1$, $\theta=16$, $\phi=0.1$
($p_{1}=0.4$, $p_{2}=0.025$)
(b) $\lambda=3$, $\theta=1/16$, $\phi=0.01$
($p_{1}=0.0025$, $p_{2}=0.04$)
Line: approximation. Dots: simulation.
Figure 20: Efficiency for LOR with group sampling
6 Conclusions
Two-stage sequential methods have been presented to estimate the RR $p_{1}/p_{2}$, the OR $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ or their logarithmic versions using independent binary observations from two populations with parameters $p_{1}$ and $p_{2}$. The estimators are unbiased; guarantee that the relative mean square error, or the mean square error for the logarithmic versions, is less than a target value $A$ irrespective of $p_{1}$ and $p_{2}$; and approximately achieve a prescribed ratio of average sample sizes when samples are taken from each population separately. The estimators can also be used with group sampling. In this case, samples are taken simultaneously from the two populations in fixed-size groups, and individual samples are extracted from those groups as needed, with a number of samples possibly left unused at the end of the process. The properties of unbiasedness and guaranteed accuracy are maintained with group sampling (and an exact sample size ratio is imposed by the sampling). Bounds and approximate expressions have been derived for the average sample sizes and the average number of sample groups, respectively. The estimation efficiency, defined in terms of the Cramér–Rao bound, has been characterized in the same way. The efficiency is generally good, both with separate sampling and with group sampling; and is close to $1$ for small $A$. Algorithms 1 and 2 specify the estimation procedure and the properties of the estimators.
$\frac{(s_{1}-1)(s_{2}-1)}{(N_{1}-1)(N_{2}-1)}$
is an unbiased estimator of $p_{1}p_{2}$, and by means of (14) a target relative error can be guaranteed. The same expressions as in RR and LRR apply for the average sample sizes, average number of groups and efficiency. In particular, the average sample sizes with separate sampling are approximately in the desired ratio, and, for $p_{1}/p_{2}$ fixed, they scale with the inverse of $\sqrt{p_{1}p_{2}}$.
Appendix A Approximation of $\operatorname{E}[|\Delta|]$
A.1 For relative risk and log relative risk
@@
$\operatorname{Var}[1/X]$ $\approx\frac{r_{1}(r_{1}+r_{2}-1)}{(r_{2}-2)(r_{2}-1)^{2}\theta^{2}}.$ (165)
[…]
@@
$\begin{split}\textstyle\operatorname{Var}\left[\Delta\right]&=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\\ &\quad+\operatorname{Var}\left[\frac{M_{2}}{l_{2}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right],\end{split}$ (166)
where (i) corresponds to the first and fourth summands, (ii) to the second and fifth, and (iii) to the third and sixth. The claim is then that (166) is well approximated by the sum of the third and sixth terms on the right-hand side. This is substantiated next, considering the small-$\phi$ regime.
The first and fourth summands in (166) are obtained from (13), (60) and (93) as
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}\right]\phi^{2}l_{1}l_{2}$ $=\frac{r_{1}(1-p_{1})}{\lambda\theta}\approx\frac{r_{1}}{\lambda\theta},$ (167)
$\operatorname{Var}\left[\frac{M_{2}}{l_{2}}\right]\phi^{2}l_{1}l_{2}$ $=r_{2}(1-p_{2})\lambda\theta\approx r_{2}\lambda\theta=(r_{1}+\mu_{1}-\mu_{2})\lambda\theta.$ (168)
Using (11), (13) and (37), and including the rounding term $\xi$ when computing $\operatorname{E}[s_{1}]$, the second and third summands in (166) can be expressed as
@@
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $=\frac{\operatorname{E}\left[s_{1}(1-p_{1})\right]}{\lambda\theta}\approx\frac{a_{1}+\xi+b_{1}\operatorname{E}[X]}{\lambda\theta},$ (169)
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $=\frac{\operatorname{Var}\left[s_{1}\right]}{\lambda\theta}\approx\frac{b_{1}^{2}\operatorname{Var}[X]}{\lambda\theta}.$ (170)
Substituting (52) and (164) into (169) and (170), and using (39), (41), (57) and (60),
@@
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\frac{1}{\lambda\theta}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\frac{1}{A}+r_{1}+\mu_{1}+\xi,$ (171)
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\lambda\theta\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)^{\!2}\frac{2r_{1}+\mu_{1}-\mu_{2}-1}{(r_{1}-2)(r_{1}+\mu_{2}-\mu_{1})}.$ (172)
By an analogous procedure, making use of (11), (13), (38) and then (39), (41), (53), (56), (60) and (165), the last two terms in (166) can be written as

### v2

@@ Appendix A Estimation procedure and properties of the estimators
Output:
  1. Estimation $\hat{\psi}$ or $\hat{\Psi}$.
Properties:
  1. Unbiased, with relative MSE (OR) or MSE (LOR) less than $A$.
  2. Average sample sizes: (148), (149); approximate ratio $\lambda$.
  3. Efficiency with element sampling: (151); approaches 11 for $A$ small.
  4. Average number of groups: (153), (154).
  5. Efficiency with group sampling: (155).
Appendix B Approximation of $\operatorname{E}[|\Delta|]$
B.1 For relative risk and log relative risk
@@
$\operatorname{Var}\left[\frac{1}{X}\right]$ $\approx\frac{r_{1}(r_{1}+r_{2}-1)}{(r_{2}-2)(r_{2}-1)^{2}\theta^{2}}.$ (163)
The value of $\operatorname{E}[|\Delta|]$ obviously depends both on the mean of $\Delta$ and on the variability of $\Delta$ with respect to its mean. In view of (92), this variability can be understood as arising from three sources: (i) the variability of $M_{1}$ and $M_{2}$ as they directly appear in that expression; (ii) the variability of $N_{1}$ and $N_{2}$ conditioned on $s_{1}$ and $s_{2}$; and (iii) the variability of $N_{1}$ and $N_{2}$ caused by variations of $s_{1}$ and $s_{2}$ (since the values of $s_{1}$ and $s_{2}$ are obtained from $X$ this is another, indirect effect of $M_{1}$ and $M_{2}$). As will be seen, the contributions of the first two sources of variability are small compared to that of the third, and can be neglected with little error. This statement can be made more precise by expressing it in terms of variance. Applying the law of total variance [5, theorem 12.2.6],
$\operatorname{Var}\left[\Delta\right]=\operatorname{E}\left[\operatorname{Var}\left[\Delta\mid M_{1},M_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\Delta\mid M_{1},M_{2}\right]\right].$ (164)
Substituting (92) into (164), and noting that $N_{1}$, $N_{2}$ depend on $M_{1}$, $M_{2}$ only through $s_{1}$ and $s_{2}$,
@@
$\begin{split}\displaystyle\operatorname{Var}\left[\Delta\right]&=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}},\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right].\end{split}$ (165)
The first three summands on the right-hand side of (165) correspond to (i), (ii) and (iii) respectively, and the fourth represents the statistical relationship between (i) and (iii) (both of which stem from the variability of $M_{1}$ and $M_{2}$).
The first summand in (165) can be written, taking into account that $M_{1}$ and $M_{2}$ are independent, as $\operatorname{Var}[M_{1}/l_{1}]+\operatorname{Var}[M_{2}/l_{2}]$, and then applying (13) gives
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\phi^{2}l_{1}l_{2}=\frac{r_{1}(1-p_{1})}{\lambda\theta}+r_{2}(1-p_{2})\lambda\theta\approx\frac{r_{1}}{\lambda\theta}+(r_{1}+\mu_{1}-\mu_{2})\lambda\theta.$ (166)
The second summand is computed as follows. Noting that $N_{1}$ and $N_{2}$ are conditionally independent given $s_{1}$ and $s_{2}$, the conditional variance $\operatorname{Var}[N_{1}/l_{1}-N_{2}/l_{2}\mid s_{1},s_{2}]$ can be expressed as $\operatorname{Var}[N_{1}/l_{1}\mid s_{1}]+\operatorname{Var}[N_{2}/l_{2}\mid s_{2}]$, and using (13) again yields
@@
$\begin{split}\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}&=\operatorname{E}\left[\frac{s_{1}(1-p_{1})}{\lambda\theta}\right]+\operatorname{E}\left[s_{2}(1-p_{2})\lambda\theta\right]\\ &\approx\frac{\operatorname{E}[s_{1}]}{\lambda\theta}+\operatorname{E}[s_{2}]\lambda\theta.\end{split}$ (167)
Then, from (37)–(42), (52), (53) and (60)–(62), and including the rounding term $\xi$ when computing $\operatorname{E}[s_{1}]$ and $\operatorname{E}[s_{2}]$,
@@
$\begin{split}\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}&\approx\frac{1}{\lambda\theta}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\lambda\theta\left(\frac{1}{A}+\mu_{2}+\xi\right)\\ &\quad+2\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right).\end{split}$ (168)
As for the third summand, from (11), (37) and (38) it can be written as
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}=\operatorname{Var}\left[\frac{s_{1}}{l_{1}p_{1}}-\frac{s_{2}}{l_{2}p_{2}}\right]\phi^{2}l_{1}l_{2}\\ \approx\frac{b_{1}^{2}\operatorname{Var}[X]}{\lambda\theta}+\lambda\theta b_{2}^{2}\operatorname{Var}\left[\frac{1}{X}\right]+2b_{1}b_{2}\operatorname{Cov}\left[X,-\frac{1}{X}\right].$ (169)
The term $\operatorname{Cov}[X,-1/X]$ is easily obtained using (52), (53) and (60):
$\operatorname{Cov}\left[X,-\frac{1}{X}\right]=-1+\operatorname{E}\left[X\right]\operatorname{E}\left[\frac{1}{X}\right]\approx\frac{2r_{1}+\mu_{1}-\mu_{2}-1}{(r_{1}-1)(r_{1}+\mu_{1}-\mu_{2}-1)}.$ (170)
Substituting (40), (42), (162), (163) and (170) into (169) and using (60)–(62),
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\phi^{2}l_{1}l_{2}\approx\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)^{\!2}(2r_{1}+\mu_{1}-\mu_{2}-1)\\ \cdot\left(\frac{\lambda\theta}{(r_{1}-2)(r_{1}+\mu_{2}-\mu_{1})}+\frac{1}{\lambda\theta r_{1}(r_{1}+\mu_{2}-\mu_{1}-2)}+\frac{2}{r_{1}(r_{1}+\mu_{2}-\mu_{1})}\right).$ (171)
Lastly, the fourth summand in (165) can be bounded using the Cauchy-Schwarz inequality [5, proposition 6.2.8]:
$\left|2\operatorname{Cov}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}},\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\right|\\ \leq 2\sqrt{\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}.$ (172)
It will be established next that (166), (168) and the right-hand side of (172) are much smaller than (171). To this end, the following observations will be useful. From (60) and (63) it stems that $r_{1}\approx r_{2}=r_{1}+\mu_{1}-\mu_{2}$ with little approximation error. In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, a relative RMSE of $10\%$ corresponds to $A=0.01$. On the other hand, $\xi=1$ is small compared to $1/A$, and so are the values of $r_{1}$ obtained from (81); namely $r_{1}\approx\sqrt{1/A}$, as can be seen from (82) or in Figure 3. In the example, $A=0.01$ gives $r_{1}=9$ for RR and $10$ for LRR.
To show that (166) is much smaller than (171), it is convenient to study the cases $\lambda\theta\approx 1$, $\lambda\theta\gg 1$ and $\lambda\theta\ll 1$ separately. For $\lambda\theta\approx 1$, the above approximations imply that (166) and (171) reduce to $2r_{1}$ and $8(1/A+r_{1})^{2}/r_{1}$ respectively, and their ratio,
$\frac{r_{1}^{2}}{4(1/A+r_{1})^{2}},$
is much smaller than $1$. For $\lambda\theta\gg 1$, (166) and (171) are approximated as $r_{1}\lambda\theta$ and $2\lambda\theta(1/A+r_{1})^{2}/r_{1}$ respectively, and their ratio is
$\frac{r_{1}^{2}}{2(1/A+r_{1})^{2}}.$
which is again small compared with 11. The case $\lambda\theta\ll 1$ gives the same result. Thus, for any value of $\lambda\theta$, the first summand in (165) can be approximately neglected in comparison with the third.
Proceeding analogously to compare (168) with (171), for $\lambda\theta\approx 1$ their approximate values are $4/A+2r_{1}$ and $8(1/A+r_{1})^{2}/r_{1}$ respectively, with a ratio
$\frac{(4/A+2r_{1})r_{1}}{8(1/A+r_{1})^{2}}=\frac{(1/A+r_{1}/2)r_{1}}{2(1/A+r_{1})^{2}}<\frac{r_{1}}{2(1/A+r_{1})},$
which is significantly smaller than $1$. For $\lambda\theta\gg 1$, (168) and (171) are approximated as $\lambda\theta/A$ and $2\lambda\theta(1/A+r_{1})^{2}/r_{1}$ respectively, and this gives a ratio
$\frac{(1/A)r_{1}}{2(1/A+r_{1})^{2}}<\frac{r_{1}}{2(1/A+r_{1})},$
which is again small compared with $1$. For $\lambda\theta\ll 1$ the result is the same. Thus (168) can also be neglected to a good approximation.
As for (172), dividing its right-hand side by the left-hand side of (171) gives twice the square root of the ratio between (166) and (171). Therefore the fourth summand in (165) is also significantly smaller than the third.
The conclusion of the preceding analysis is that (165) can be approximated by keeping only the third summand in the right-hand side, as it is significantly larger than the others. This means that the variability in $\Delta$ is mostly due to the variability of $N_{1}$ and $N_{2}$ caused by variations of $s_{1}$ and $s_{2}$, i.e. (iii) as defined above. The variability of $M_{1}$ and $M_{2}$, (i), and the variability of $N_{1}$ and $N_{2}$ conditioned on $s_{1}$ and $s_{2}$, (ii), are comparatively smaller. Therefore, to compute $\operatorname{E}[|\Delta|]$, the variables $M_{1}$, $M_{2}$ in (92) can be replaced by their means, and $N_{1}$, $N_{2}$ can be replaced by their conditional means given $s_{1}$, $s_{2}$, which yields (94).
B.2 For odds ratio and log odds ratio
The law of total variance [5, theorem 12.2.6] conditioning on $\bar{M}_{1}$ and $\bar{M}_{2}$ gives, for OR and LOR,
$\operatorname{Var}\left[\Delta\right]=\operatorname{E}\left[\operatorname{Var}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right].$ (173)
The variables $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$, and therefore $N_{1}$, $N_{2}$, depend on $M_{1}$, $M_{2}$ only through $\bar{M}_{1}$, $\bar{M}_{2}$. This implies that $M_{1}$, $M_{2}$, $N_{1}$, $N_{2}$, are conditionally independent given $\bar{M}_{1}$, $\bar{M}_{2}$. Thus, using (92), the first term in (173) is written as
$\operatorname{E}\left[\operatorname{Var}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right]=\operatorname{E}\left[\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right],$ (174)
whereas the second term is